OpenStudy (anonymous):

5x+4y=20 graph equation

OpenStudy (kc_kennylau):

https://www.desmos.com/calculator/b8cvnrjjze

OpenStudy (a1234):

Put it in slope-intercept form: 5x + 4y = 20 4y = -5x + 20 y = -5/4x + 5

OpenStudy (a1234):

Then graph it.